What is Joule - Thomson effect ? how is Joule - Thomson coefficient determined experimentally ?
1
Expert's answer
2019-11-15T07:37:55-0500

Solution.

The Joule-Thomson effect is a phenomenon in which there is a change in the temperature of a gas or liquid under stationary adiabatic throttling. This process occurs when the gas flows slowly under the influence of a constant pressure drop through the porous partition.

Expression for Joule-Thomson coefficient:

"\mu = \frac{dT}{dP} = \frac{T(\frac{\delta V}{\delta T})p - V}{Cp}"

For an ideal gas, this expression is 0, and for a real gas, it is determined by the equation of state.

Answer:

The Joule-Thomson effect is a phenomenon in which there is a change in the temperature of a gas or liquid under stationary adiabatic throttling. This process occurs when the gas flows slowly under the influence of a constant pressure drop through the porous partition.

Expression for Joule-Thomson coefficient:

"\mu = \frac{dT}{dP} = \frac{T(\frac{\delta V}{\delta T})p - V}{Cp}"

For an ideal gas, this expression is 0, and for a real gas, it is determined by the equation of state.